# Phonon-induced renormalization of exchange interactions in metallic 2D magnets

### v1

@@ Abstract
Introduction.— Magnetism is a quantum phenomenon that plays an important role in technology. Although fundamental aspects of magnetism had been established decades ago, understanding the behavior of magnetic materials at the microscopic scale is still the subject of intensive research. Recently, the interest to fundamental aspects of magnetism has been revived due to the emergence of novel materials that host low-dimensional forms of magnetism, allowing for its efficient manipulation and control Gong et al. (2017); Huang et al. (2017); Klein et al. (2018); Huang et al. (2018); Jiang et al. (2018). Among these materials, metallic 2D magnets with the most prominent example of Fe3GeTe2 Deng et al. (2018); Fei et al. (2018) constitutes a special class of objects where charge carriers provide additional degree of freedom for tailoring the magnetic properties.
@@
where $G_{ij}^{\sigma}(i\omega_{n})$ is the spin-polarized electron propagator, $\Delta_{i}$ is the exchange splitting of electrons at lattice site $i$, and $\mathrm{Tr}_{\omega L}$ denote the sum over Matsubara frequencies $i\omega_{n}$ and orbital indices $L$. In most of the practical situations, $\Delta_{i}$ and $G_{ij}^{\sigma}$ are derived from single-particle Hamiltonians, disregarding explicit treatment of many-body effects. A modification of Eq. (2) has been proposed in the context of strongly correlated systems Katsnelson and Lichtenstein (2000), giving the possibility to obtain low-temperature corrections to $J_{ij}$ due to renormalization of electron-electron interactions; in this case $\Delta$ is replaced by spin-dependent part of the local electron self-energy. In some cases, such renormalization even results in a sign reversal $J_{ij}$, as demonstrates the example of Mott insulators in the presence of strong laser field Mentink et al. (2015); Itin and Katsnelson (2015); Mikhaylovskiy et al. (2015). At nonzero temperatures, thermal lattice fluctuations are supposed to provide another important effect on the exchange interactions, especially at low dimensions where electronic structure is more susceptible to perturbations of the atomic structure. Earlier attempts to describe spin-lattice effects in 2D magnets were limited to insulators, such as CrI3 Sadhukhan et al. (2022), with the main effect originating from a direct modification of the hopping integrals upon atomic displacements. In contrast, metallic magnets allow for a different mechanism of spin-lattice interaction related to the coupling of phonons with conducting electrons that is strongly temperature dependent.
@@
$+\sum_{\bf q}\omega_{\bf q}b^{{\dagger}}_{\bf q}b_{\bf q}+\sum_{{\bf q},<ij>\sigma}g_{\bf q}(b_{\bf q}^{{\dagger}}+b_{-{\bf q}})c_{i\sigma}^{{\dagger}}c_{j\sigma},$ (4)
@@
$\Sigma^{\sigma}_{\mathbf{k}}(\omega,T)=2\lambda\frac{k_{B}T}{N^{\sigma}_{F}}\sum_{\mathbf{q}}G^{\sigma}_{\mathbf{k+q}}(\omega),$ (5)
@@
where $c$ is a renormalization constant determined by the electronic structure of the system. This expression assumes that the electron-phonon contribution to the self-energy can be considered as a perturbation and taken into account in the linear approximation which seems to be reasonable in the most of cases. Importantly, the temperature correction $\delta J=-c\lambda T$ does not scale with $J(0)$ meaning that the renormalization effect is more significant when $J(0)\lesssim c\lambda T$. As we are mostly interested in the ordered phase below the critical temperature, one can recast the renormalization criterion as $c\gtrsim T_{\mathrm{c}}/\lambda T$. In Fig. 1, we provide these constants fitted for the three cases considered. We emphasize that the renormalization effect persists even for spin-independent electron-phonon coupling, i.e., when $\lambda^{\uparrow}=\lambda^{\downarrow}$. For unequal $\lambda^{\uparrow}$ and $\lambda^{\downarrow}$ the effect is expected to be stronger.
@@
Figure 2: (Top) Spin-resolved electron spectral function of monolayer Fe3GeTe2 calculated in the presence of the electron-phonon coupling at $T=100$ K for the states near the Fermi level. Original single-particle DFT band structure ($T=0$) is shown by the red solid line. (Bottom) Temperature dependence of the renormalized leading exchange interactions shown schematically in the inset. The numerical values of $J_{ij}$ are given in the Supplemental Material SM .
@@
The electronic structure of Fe3GeTe2 in the vicinity of the Fermi energy is mainly formed by Fe $(3d)$ hybridized with Te $(5p)$ states Badrtdinov et al. (2023), which allows us to parameterize them by maximally localized Wannier functions Marzari and Vanderbilt (1997); Marzari et al. (2012). Using such parametrization, we construct a tight-binding Hamiltonian in the basis of Wannier functions to calculate the electron propagators as well as electron-phonon matrix elements needed to obtain the electron self-energy in the Migdal approximation (see the SM for details SM ). For this purpose, we use the epw code Poncé et al. (2016); Giustino (2017) modified to treat spin-polarized states. Figure 2 shows spin-resolved spectral function calculated for monolayer Fe3GeTe2 at $T=0$ and $T=100$ K. One can see that spin-down electrons are more susceptible to the electron-phonon coupling, which is consistent with distinct coupling constants $\lambda^{\uparrow}\sim$ 0.5 and $\lambda^{\downarrow}\sim$ 0.26 Badrtdinov et al. (2023). Such behavior indicates a phonon-induced renormalization of the exchange splitting $\Delta_{\bf k}(\omega)$, which provides a direct contribution to the exchange interactions $J_{ij}$ according to Eq. (2).
The resulting temperature dependence of the renormalized leading exchange interactions between iron atoms in Fe3GeTe2 is presented in Fig. 2 (see the SM for further details SM ). Overall, the effect of exchange renormalization is not large, and found to be within 10% for $T=200$ K. This can be ascribed to a relatively large effective $\tilde{t}/\tilde{\Delta}$ ratio, in accordance with the results presented in Fig. 1. Indeed, our first-principles calculations yield the average exchange splitting $\tilde{\Delta}$ of Fe($d$) electrons to be $\sim$2 eV, while the effective hopping $\tilde{t}$ determined from the bandwidth $W_{d}\simeq$ 10 eV is estimated to be $\tilde{t}\simeq W_{d}/2N_{nn}\approx$ 1.6 eV, where $N_{nn}=3$ is the number of nearest neighbors in the honeycomb lattice. The resulting effective parameters ratio $\tilde{t}/\bar{\Delta}\sim$ 0.8 suggests a small renormalization effect, in agreement with Fig. 1. The corresponding renormalization constant $c$ introduced in Eq. (6) is estimated to be around 0.1–0.2 in Fe3GeTe2, which also indicates weak renormalization. Despite the effect is relatively small, the renormalization can still affect the observables such as the spectrum of spin-wave excitations or the Curie temperature.
@@
where $J_{\mu\nu}({\bf q})$ are the Fourier transform of the exchange interaction matrix, $\Phi=1-\left[1-\braket{S^{2}_{z}}/2\right]$ is the Anderson-Callen decoupling factor Anderson and Callen (1964); Fröbrich et al. (2000) for $S=1$, which satisfies the kinematic condition. The indices $\mu$ and $\nu$ run from 1 to 3 over magnetic sublattices.
Figure 3 shows the calculated spin-wave spectra that exhibit two optical branches and one acoustic branch dispersing quadratically $\Omega_{\bf q}\simeq\Omega_{0}+Dq^{2}$ near the $\Gamma$ point, where $D$ is the spin-stiffness constant and $\Omega_{0}$ is the SIA-induced gap of the order of meV. Temperature dependence of the magnetization $\braket{S_{z}}$ already leads to a softening of the magnon branches, which drops to 0 at $T=T_{\mathrm{C}}$. As the same time, taking into account the temperature dependence of magnetic interactions increases this softening even stronger, reducing the spin-stiffness constant $D$.

### v2

# Phonon-induced renormalization of exchange interactions in metallic two-dimensional magnets

@@ Abstract
Introduction.— Magnetism is a quantum phenomenon that plays an important role in technology. Although fundamental aspects of magnetism had been established decades ago, understanding the behavior of magnetic materials at the microscopic scale is still the subject of intensive research. Recently, the interest to fundamental aspects of magnetism has been revived due to the emergence of novel materials that host low-dimensional forms of magnetism, allowing for its efficient manipulation and control Gong et al. (2017); Huang et al. (2017); Klein et al. (2018); Huang et al. (2018); Jiang et al. (2018). Among these materials, metallic two-dimensional (2D) magnets with the most prominent example of Fe3GeTe2 Deng et al. (2018); Fei et al. (2018) constitutes a special class of objects where charge carriers provide additional degree of freedom for tailoring the magnetic properties.
@@
where $G_{ij}^{\sigma}(i\omega_{n})$ is the spin-polarized electron propagator, $\Delta_{i}$ is the exchange splitting of electrons at lattice site $i$, and $\mathrm{Tr}_{\omega L}$ denote the sum over Matsubara frequencies $i\omega_{n}$ and orbital indices $L$. In most of the practical situations, $\Delta_{i}$ and $G_{ij}^{\sigma}$ are derived from single-particle Hamiltonians, disregarding explicit treatment of many-body effects. A modification of Eq. (2) has been proposed in the context of strongly correlated systems Katsnelson and Lichtenstein (2000), giving the possibility to take into account renormalization of $J_{ij}$ due to electron-electron interactions. In this case $\Delta$ is replaced by spin-dependent part of the local electron self-energy $\Sigma^{\sigma}(i\omega_{n})$, obtained self-consistently within dynamical mean-field theory (DMFT). This self-energy accounts for dynamical spin fluctuations and ensures temperature-dependence of the exchange.
An alternative way to describe temperature dependence of $J_{ij}$ due to magnetic fluctuations is to introduce magnetic disorder within the formalism of disordered local moments Pindor et al. (1983); Staunton et al. (1985, 1986). The degree of disorder is temperature dependent and typically determined from atomistic spin simulations for each temperature. Subsequently, Eq. (2) is modified to take the disorder into account Böttcher et al. (2012); Szilva et al. (2013). This approach can be considered as a semiclassical approximation to DMFT Niklasson et al. (2003). Recently, semianalytical methods have been proposed to treat this problem, in which temperature-dependent corrections to the exchange are evaluated from spin correlation functions Rózsa and Atxitia (2023).
At nonzero temperatures, thermal lattice fluctuations provide additional contribution to the temperature dependence of exchange interactions Mankovsky et al. (2020). This effect is supposed to be more pronounced in low dimensions where electronic structure is more susceptible to perturbations of the atomic structure. Earlier attempts to describe spin-lattice effects in 2D magnets were limited to insulators, such as CrI3 Sadhukhan et al. (2022), with the main effect originating from a direct modification of the hopping integrals upon atomic displacements. In contrast, metallic magnets allow for a different mechanism of spin-lattice interaction related to the coupling of phonons with conducting electrons that is strongly temperature dependent.
@@
$+\sum_{\bf q}\omega_{\bf q}b^{{\dagger}}_{\bf q}b_{\bf q}+\sum_{{\bf q},\langle ij\rangle\sigma}g_{\bf q}(b_{\bf q}^{{\dagger}}+b_{-{\bf q}})c_{i\sigma}^{{\dagger}}c_{j\sigma},$ (4)
@@
$\Sigma^{\sigma}_{\mathbf{k}}(\omega,T)\simeq 2\lambda\frac{k_{B}T}{N^{\sigma}_{F}}\sum_{\mathbf{q}}G^{\sigma}_{\mathbf{k+q}}(\omega),$ (5)
@@
where $c$ is a renormalization constant determined by the electronic structure of the system. This expression assumes that the electron-phonon contribution to the self-energy can be considered as a perturbation and taken into account in the linear approximation which seems to be reasonable in the most of cases. Importantly, the temperature correction $\delta J=-c\lambda T$ does not scale with $J(0)$ meaning that the renormalization effect is more significant when $J(0)\lesssim c\lambda T$. As we are mostly interested in the ordered phase below the critical temperature, one can recast the renormalization criterion as $c\gtrsim T_{\mathrm{c}}/\lambda T$, keeping in mind the estimate $T_{\mathrm{c}}\sim J(0)$. In Fig. 1, we provide these constants fitted for the three cases considered. We emphasize that the renormalization effect persists even for spin-independent electron-phonon coupling, i.e., when $\lambda^{\uparrow}=\lambda^{\downarrow}$. For unequal $\lambda^{\uparrow}$ and $\lambda^{\downarrow}$ the effect is expected to be stronger.
@@
Figure 2: (Top) Spin-resolved electron spectral function of monolayer Fe3GeTe2 calculated in the presence of the electron-phonon coupling at $T=100$ K for the states near the Fermi level. Left and right panels correspond to the spin-up and spin-down states, respectively. Original single-particle DFT band structure ($T=0$) is shown by the red solid line. (Bottom) Temperature dependence of the renormalized leading exchange interactions shown schematically in the inset. The numerical values of $J_{ij}$ are given in the Supplemental Material SM (see also references Liechtenstein et al. (1987); Mazurenko and Anisimov (2005); Tyablikov (1983); Nolting and Ramakanth (2009); Fröbrich and Kuntz (2006) therein).
@@
The electronic structure of Fe3GeTe2 in the vicinity of the Fermi energy is mainly formed by Fe $(3d)$ hybridized with Te $(5p)$ states Badrtdinov et al. (2023), which allows us to parameterize them by maximally localized Wannier functions Marzari and Vanderbilt (1997); Marzari et al. (2012). Using such parametrization, we construct a tight-binding Hamiltonian in the basis of Wannier functions to calculate the electron propagators as well as electron-phonon matrix elements needed to obtain the electron self-energy in the Migdal approximation (see the SM for details SM ). For this purpose, we use the epw code Poncé et al. (2016); Giustino (2017) modified to treat spin-polarized states. Fig. 2 shows spin-resolved spectral function calculated for monolayer Fe3GeTe2 at $T=0$ and $T=100$ K. One can see that spin-down electrons are more susceptible to the electron-phonon coupling, which is consistent with distinct coupling constants $\lambda^{\uparrow}\sim$ 0.5 and $\lambda^{\downarrow}\sim$ 0.26 Badrtdinov et al. (2023). Such behavior indicates a phonon-induced renormalization of the exchange splitting $\Delta_{\bf k}(\omega)$, which provides a direct contribution to the exchange interactions $J_{ij}$ according to Eq. (2).
The resulting temperature dependence of the renormalized leading exchange interactions between iron atoms in Fe3GeTe2 is presented in Fig. 2 (see the SM for further details SM ). Overall, the effect of exchange renormalization is not large, and found to be within 10% for $T=200$ K. This can be ascribed to a relatively large effective $\tilde{t}/\tilde{\Delta}$ ratio, in accordance with the results presented in Fig. 1. Indeed, our first-principles calculations yield the average exchange splitting $\tilde{\Delta}$ of Fe($d$) electrons to be $\sim$2 eV, while the effective hopping $\tilde{t}$ determined from the bandwidth $W_{d}\simeq$ 10 eV is estimated to be $\tilde{t}\simeq W_{d}/2N_{nn}\approx$ 1.6 eV, where $N_{nn}=3$ is the number of nearest neighbors in the honeycomb lattice. The resulting effective parameters ratio $\tilde{t}/\bar{\Delta}\sim$ 0.8 suggests a small renormalization effect, in agreement with Fig. 1. The corresponding renormalization constant $c$ introduced in Eq. (6) is estimated to be around 0.1–0.2 in Fe3GeTe2, which also indicates weak renormalization. Despite the effect is relatively small, the renormalization can still affect the observables such as the spectrum of spin-wave excitations or the Curie temperature. We note that Eq. (2) for the exchange is only applicable to the ordered phase ($T<T_{\mathrm{C}}$), since in the paramagnetic phase both $\Delta_{i}$ and $\braket{S_{i}^{z}}$ tends to zero, and this limit requires a special careful treatment Pourovskii (2016); Katanin et al. (2023). Here we restrict ourselves only to the ferromagnetic phase leaving the problem of electron-phonon interaction in the paramagnetic phase for the future studies.
@@
where $J_{\mu\nu}({\bf q})$ are the Fourier transform of the exchange interaction matrix, $\Phi=1-\left[1-\braket{S^{2}_{z}}/2\right]$ is the Anderson-Callen decoupling factor Anderson and Callen (1964); Fröbrich et al. (2000) for $S=1$, which satisfies the kinematic condition. The indices $\mu$ and $\nu$ run from 1 to 3 over magnetic sublattices. From Eq. (8) follows that the magnon energies $\Omega_{\bf q}\sim\braket{S_{z}}$, which is not always correct because the magnon excitations are known to exist beyond the critical temperature, assuming strong short-range magnetic order. Here, we are not interested in the paramagnetic regime, but note that the behavior at $T>T_{\mathrm{C}}$ can be restored by introducing short-range spin-order parameter as it is done, e.g., in self-consistent spin-wave theory Irkhin et al. (1999).
Fig. 3 shows the calculated spin-wave spectra of monolayer Fe3GeTe2 that exhibit two optical branches and one acoustic branch dispersing quadratically $\Omega_{\bf q}\simeq\Omega_{0}+Dq^{2}$ near the $\Gamma$ point, where $D$ is the spin-stiffness constant and $\Omega_{0}$ is the SIA-induced gap of the order of meV. Temperature dependence of the magnetization $\braket{S_{z}}$ already leads to a softening of the magnon branches, which drops to 0 at $T=T_{\mathrm{C}}$. As the same time, taking into account the temperature dependence of magnetic interactions increases this softening even stronger, reducing the spin-stiffness constant $D$.
@@
Supplemental Materials: Phonon-induced renormalization of exchange interactions in metallic two-dimensional magnets
Temperature-dependent exchange interactions were calculated using local force theorem approach S_Ref1 ; S_Ref2 :
[…]
Here $\widetilde{\Delta}_{i}(\omega,T)$ and $\widetilde{G}^{\sigma}_{ij}(\omega,T)$ stands for Fourier transformed component of correlated intra-orbital spin-splitting energy and Green’s function due to electron-phonon coupling:
$\widetilde{\Delta}_{i}(\omega,\mathbf{k},T)=H_{i}^{\uparrow}(\mathbf{k})-H_{i}^{\downarrow}(\mathbf{k})+\Sigma^{\uparrow}(\omega,\mathbf{k},T)-\Sigma^{\downarrow}(\omega,\mathbf{k},T),$ (S2)
$\widetilde{G}_{ij}^{-1}(\omega,\mathbf{k},T)=G_{ij}^{-1}(\omega,\mathbf{k})-\Sigma(\omega,\mathbf{k},T).$
The electron self-energy of electron-phonon interaction in Migdal approximation has the form
$\Sigma^{\sigma}_{n\mathbf{k}}(\omega,T)=\sum_{m\,\mathbf{q}\nu}|g^{\sigma}_{mn,\nu}(\mathbf{k},\mathbf{q})|^{2}\times\left[\frac{b_{\mathbf{q}\nu}+f^{\sigma}_{m\mathbf{k+q}}}{\omega-\varepsilon^{\sigma}_{m\mathbf{k+q}}+\hbar\omega_{\mathbf{q\nu}}+i\eta}+\frac{b_{\mathbf{q}\nu}+1-f^{\sigma}_{m\mathbf{k+q}}}{\omega-\varepsilon^{\sigma}_{m\mathbf{k+q}}-\hbar\omega_{\mathbf{q\nu}}+i\eta}\right],$ (S3)
where $b_{\mathbf{q}\nu}=(\exp[\hbar\omega_{\mathbf{q\nu}}/k_{B}T]-1)^{-1}$ corresponds to the Bose occupation function for phonons with wave vector $\mathbf{q}$, mode index $\nu$ and frequency $\omega_{\mathbf{q\nu}}$. In turn, $f^{\sigma}_{m\mathbf{k}}=(\exp[(\varepsilon^{\sigma}_{m\mathbf{k}})/k_{B}T]+1)^{-1}$ is the Fermi occupation function for the electron states with energy $\varepsilon^{\sigma}_{m\mathbf{k}}$ given with respect to the Fermi level. The electron-phonon matrix elements contain the information about the derivative of self-consisted spin dependent electronic potential $\partial_{\mathbf{q}\nu}V^{\sigma}$ in the basis of Bloch functions $\psi^{\sigma}_{n\mathbf{k}}$:
$g^{\sigma}_{mn,\nu}(\mathbf{k},\mathbf{q})=\sqrt{\frac{\hbar}{2M\omega_{\mathbf{q\nu}}}}\braket{\psi^{\sigma}_{m\mathbf{k+q}}}{\partial_{\mathbf{q}\nu}V^{\sigma}}{\psi^{\sigma}_{n\mathbf{k}}}.$ (S4)
Expectation values of $\braket{S_{z}}$ and $\braket{S^{2}_{z}}$ in case of $S=1$ are the following S_Ref3 ; S_Ref4 ; S_Ref5 :
$\braket{S_{z}}=S\frac{1+2\sum_{\mathbf{q}\nu}b_{\mathbf{q}\nu}}{1+3\sum_{\mathbf{q}\nu}b_{\mathbf{q}\nu}+3(\sum_{\mathbf{q}\nu}b_{\mathbf{q}\nu})^{2}},$ (S5)
$\braket{S^{2}_{z}}=\frac{2\sum_{\mathbf{q}\nu}b_{\mathbf{q}\nu}}{1+3\sum_{\mathbf{q}\nu}b_{\mathbf{q}\nu}}+\frac{1+\sum_{\mathbf{q}\nu}b_{\mathbf{q}\nu}}{1+3\sum_{\mathbf{q}\nu}b_{\mathbf{q}\nu}}\braket{S_{z}}$
Table 1: Temperature evolution of exchange interaction (in meV) of Fe3GeTe2 due to electron-phonon coupling.
$i$ $d(\AA)$ $\mathrm{Fe_{i}-Fe_{j}}$ $J$(T = 0 K) $J$(T = 50 K) $J$(T = 100 K) $J$(T = 200 K) $J$(T = 300 K) $J(300)/J(0)$
1 2.493 I - I -110.1 -107.4 -106.8 -105.0 -102.8 0.93
2 2.624 I - II -37.1 -36.5 -36.2 -35.8 -35.2 0.95
3 3.991 (I - I)/(II - II) 1.6/7.7 1.6/7.3 1.6/7.2 -1.6/6.9 -1.6/6.6 1.0/0.85
4 4.712 I - I 1.4 1.2 1.1 1.0 0.9 0.64
5 4.782 I - II -6.4 -6.2 -6.1 -5.9 -5.6 0.88
6 6.234 I - II 1.6 1.6 1.6 1.5 1.4 0.88
7 6.926 (I - I)/(II - II) -0.8/-3.6 -1.0/-3.4 -0.8/-3.2 -0.7/-3.0 -0.6/-2.8 0.75/0.77
8 7.361 I - I -2.0 -1.8 -1.7 -1.7 -1.6 0.80
9 7.998 (I - I)/(II - II) -2.0/0.5 -2.0/0.5 -2.0/0.4 -1.9/0.4 -1.9/0.3 0.95/0.60
10 8.377 I - I -0.8 -0.8 -0.8 -0.8 -0.7 0.87